OpenStudy (anonymous):

Solve the triangle. Round lengths to the nearest tenth and angle measures to the nearest degree. a = 7, b = 7, c = 5

OpenStudy (wolf1728):

ARRGGHH!!!!!!!!!!!! Had it al ready, page faded then came back with nothing Attached graphic still there

OpenStudy (wolf1728):

cosine Angle A = 2.5 / 7 cosine Angle A = 0.3571428571 Angle A = Arc cos (0.3571428571) = 69.075 Angle B = Angle B = 69.075 Angle C = 180 - 2*69.075 Angle C = 41.85

OpenStudy (wolf1728):

Hope that helped :-)

OpenStudy (anonymous):

Thank you! :)